# Mixed Strategy Constraints in Continuous Games

### v1

@@ Abstract.
Equilibrium problems representing interaction in physical environments typically require continuous strategies which satisfy opponent-dependent constraints, such as those modeling collision avoidance. However, as with finite games, mixed strategies are often desired, both from an equilibrium existence perspective as well as a competitive perspective. To that end, this work investigates a chance-constraint-based approach to coupled constraints in generalized Nash equilibrium problems which are solved over pure strategies and mixing weights simultaneously. We motivate these constraints in a discrete setting, placing them on tensor games ($n$-player bimatrix games) as a justifiable approach to handling the probabilistic nature of mixing. Then, we describe a numerical solution method for these chance constrained tensor games with simultaneous pure strategy optimization. Finally, using a modified pursuit-evasion game as a motivating examples, we demonstrate the actual behavior of this solution method in terms of its fidelity, parameter sensitivity, and efficiency.
conference: ; ;
@@ 1. Introduction
When solving Nash equilibrium problems for applications in robotics, control, and automation, there an assortment of practical considerations to be addressed. These games are typically played in physical space, and therefore are modeled with continuous strategy spaces. As with finite games, mixed strategies are often required for continuous games to ensure the existence of an equilibrium. However, physical-space games often also have constraints involving the strategies of multiple players (e.g. collision constraints: no two players can occupy the same space at the same time). Properly handling constraints in a mixed-strategy framework is the focus of this work.
A more simple application of these so-called coupling constraints occurs in discrete games with mixed strategies, before any explicit consideration of improving the pure strategies. Coupling constraints are troublesome in the presence of mixing, even in discrete settings: since mixed strategies are probabilistic, a constraint may or may not be violated depending on the strategies actually realized by opponents. We determine that chance constraints can be employed to intuitively account for mixing.
Of greater interest, however, are continuous-strategy games. In general, continuous games require mixing over an infinite support of pure strategies to ensure the existence of a Nash equilibrium. However, in certain classes of games such as separable games, finite support mixed strategy equilibria exist ((Stein et al., 2008)). We seek pure strategies that admit a competitive mix with low cost, the pure strategy space can itself be subject to optimization, as in (Peters et al., 2022). In practice, mixing is quite desirable in continuous games, often providing a competitive advantage even with minimal support.
In this paper, we place multiple-player constraints on mixed strategy games subject to continuous pure strategy optimization. We employ chance constraints in this continuous setting, describing in detail the necessary modifications required to function in a pure strategy optimization setting. Keeping in mind our goal of formulating games in physical space, we consider a modified, three-player constrained pursuit-evasion game. We conclude with promising areas of further investigation, and a discussion of the application of these constraints to model common problems in our spatial domain.
@@ 1.1. Our Contributions
  • A formulation of chance constraints intended to be applied to mixed-strategy tensor games with tensor constraints, a bimatrix game generalization;
  • A numerical solver which uses this formulation to seek mixed-strategy Nash equilibria, optimizing along both the pure strategy space and the mixing weights;
  • Empirical behavior of the solver when applied to constrained pursuit-evasion games,
  • Analysis of the solver quality with respect to its parameters, particularly regarding the feasibility of the resulting solution.
The structure of this paper is as follows: In the remainder of this section we briefly discuss the related work on generalized Nash equilibrium problems and chance constraints to scope our work. In Section 2 we define tensor games with tensor constraints, a mathematical preliminary to working with chance constraints under pure strategy optimization, and in Section 3 we discuss those constraints in discrete settings. Section 4 covers our solver with which we compute the empirical results discussed in Section 5, followed by a brief consideration of future extensions and conclusion.
@@ 1.2.1. Coupled constraints
For our purposes in constraining game-theoretic behavior in real space, we require a notion of constraints involving multiple players. These constraints are known as shared or coupled constraints, and games [^1] There is some nuance as to whether GNEPs are, in fact, classifiable as games. In physical-space problems where mutual interactions are necessary and their corresponding constraints are as natural as single-player constraints, this distinction is of little importance. which feature them are known as generalized Nash equilibrium problems, or GNEPs, as described in (Facchinei and Kanzow, 2010a). Other (older) terminology, like that seen in (Harker, 1991), refers to coupled constraints as feasible set interactions, found in social equilibrium problems and pseudo-Nash games.
Solutions to solving GNEPs vary, and generally do not use an explicit concept of mixing. (Krawczyk, 2007) summarizes numerical solutions to solving these games, a project further advanced in, e.g., (von Heusinger, 2009). Past avenues include penalty methods (Facchinei and Kanzow, 2010b), augmented Lagrangian methods (Kanzow and Steck, 2016), and methods informed by error bound theory (Ba and Pang, 2022). This work focuses on the representation and modeling of constraints in mixed-strategy games, and less on computation. In this work we characterize GNEPs as a mixed complementarity problem via the KKT conditions, and solve them via a readily available solver (Dirkse and Ferris, 1995).
Our primary divergence from these works lies in the presence of pure strategy optimization, where we consider both pure strategies and their mixing weights in finding a Nash equilibrium. To our knowledge, this specific form of generalized Nash equilibrium has never been fully explored in the literature, which prevents further development of mixed-strategy game-theoretic agents in constrained physical-space games.
@@ 1.2.2. Chance constraints
To intuitively account for the stochasticity of mixed strategies, we use chance constraints, originally formulated in (Charnes and Cooper, 1959). Chance-constrained problems without feasible set interactions (that is, in ordinary Nash equilibrium problems) are well covered, and significant work has been done to characterize the circumstances in which such problems have equilibria (as in (Henrion and Strugarek, 2008), (Geletu et al., 2013)). Chance constraints have been used in the presence of mixing for decades, see e.g. (Mukherjee, 1980).
Some work has been done describing the behavior of chance constraints in GNEPs, but never with the intent of pure strategy optimization. (Zhang et al., 2022) manage mixing chance constraints in discrete mixed-strategy scenarios from the variational inequality point of view, while (Singh and Lisser, 2018) apply the same treatment to a continuous single strategy to handle stochastic payoffs, but the former assumes given pure strategies, while the latter does not mix. We assume pure strategies as additional, continuous decision variables rather than as given, and mix over them under the presence of chance constraints.
2. Preliminaries: Tensor Games with Tensor Constraints
We begin with then necessary mathematical preliminaries to discuss (generalized) Nash equilibrium problems under pure strategy optimization when all players interact in constraints.
In an $n$-player game, the resulting expected cost is a tensor, with each player mixing along each axis of the tensor. As such we consider tensor games, (sometimes referred to as multilinear games), $N$-player extensions of bimatrix games. Such games are established in the literature ((Papadimitriou and Roughgarden, 2005), (Huang and Qi, 2017)); they are known to have mixed-strategy equilibria and multiple solution methods exist. We specifically formulate tensor games with tensor constraints, allowing all players to affect all constraints through their mixing probabilities or, in continuous games, through the cost and constraint tensors.
2.1. Definitions
Tensor games occur when each player $i$ mixes over $m_{i}$ pure strategies with probabilities $x_{i}$. We denote player $i$’s $k$-th pure strategy $s_{i,k}$ and that player’s pure strategy space as $\mathcal{S}_{i}$, such that $s_{i}\in\mathcal{S}_{i}^{m_{i}}$ (and similarly denoting $s\in\mathcal{S}$).
In tensor games, each player has a cost tensor; each element of the cost tensor corresponds to a joint strategy. Cost tensors are denoted $A_{i}$, and all $A_{i}\in\mathbb{R}^{m_{1},..,m_{i}}$. The resulting cost, when each player has probability $x_{i,k}$ assigned to their $k$th strategy, is a homogeneous $N$-polynomial. For ease, we denote the operation that forms this polynomial $A_{i}\;[\cdot]\;x$, under the following definition:
Definition 2.1 ().
@@ 2.1. Definitions
(1) $T\;[\cdot]\;x:=(((T\cdot x_{i})\cdot..)\cdot x_{N}$
We also constrain tensor games with a total of $C$ constraints (indexed by $j)$. Each player may or may not respect a particular constraint; we denote the index set of constraints respected by player $i$ as $\mathcal{O}_{i}$. Each constraint tensor is denoted as $Q_{j}$ and has the same shape as the cost tensor. The elements of the constraint tensor represent some value regarding the feasibility of the joint strategy with the corresponding index, and we constrain the expected value: $Q_{j}\;[\cdot]\;x\geq\epsilon_{j}$ for some arbitrary $\epsilon_{j}$.
With this notation in place, we define tensor games with tensor constraints:
Definition 2.2 ().
@@
(2) $\forall i\in\{1..N\}:\begin{bmatrix}\;\;\min\limits_{x_{i}\in\Delta^{m_{i-1}}}\;\;A_{i}\;[\cdot]\;x\;\;\;\;\textrm{s.t.}\;\;\;\;\\[8.0pt] \;\;\;\;\;\;Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}\geq 0\;\;\forall j\in\mathcal{O}_{i}\end{bmatrix}$
Let $x_{\neg i}$ denote the concatenation of strategy vectors $x_{1},...,x_{i-1},x_{i+1},...,x_{N}$, and let $(x_{i},x_{\neg i})$ be shorthand for $x$.
Definition 2.3 ().
An equilibrium point for a game $G(A,Q,\epsilon)$ is a strategy vector $x^{*}$ such that for all $i\in\{1,...,N\}$,
(3) $A_{i}\;[\cdot]\;(x_{i}^{*},x_{\neg i}^{*})\leq A_{i}\;[\cdot]\;(x_{i},x_{\neg i}^{*})\ \forall x_{i}\in\Delta^{m_{i-1}}:Q_{i,j}\;[\cdot]\;(x_{i},x_{\neg i}^{*})-\epsilon_{i,j}\geq 0\ \forall j\in\mathcal{O}_{i}$
In discrete games, the way in which the contents of the cost and constraint tensors are derived is external to the game. However, for our eventual purpose of applying constraints on continuous mixed-strategy games to model physical-space interactions, we assume that the elements of the tensors are related the pure strategies by way of real-valued cost and constraint functions $f_{i}$ and $g_{i,j}:\mathcal{S}\rightarrow\mathbb{R}$. Let $K$ be an index into a cost or constraint tensor (and therefore $K_{i}$ is the index of the strategy realized by player $i$ in a particular joint strategy). Then, the cost and constraint tensor elements are defined accordingly:
(4) $A_{i,K}=f_{i}(s_{1,K_{1}}\;..\;s_{N,K_{N}})\;\;;\;\;Q_{i,j,K}=\rho(g_{i,j}(s_{1,K_{1}}\;..\;s_{N,K_{N}}))$
@@
When using pure strategy optimization, our goal (from the perspective of player $i$) is to both mix over pure strategies correctly, and to select $s_{i}$ that generate the most favorable $A$ and $Q$ under a correct mix. For this purpose we assume $f$ and $g$ are twice differentiable and defined over a compact $\mathcal{S}$.
@@ 2.2. Existence of equilibria
Tensor games with tensor constraints do not always admit and equilibrium point. This can cause problems for the optimization of pure strategies. As such, we briefly pause to establish sufficient conditions to ensure the presence of an equilibrium in a tensor-constrained tensor game.
Theorem 2.4 ().
@@
We then follow the conditions put forth by (Dutang, 2013):
@@
    First, $Q_{j}\;[\cdot]\;x$ is continuous and concave in $x_{i}$ given all $x_{\neg i}$, since it is linear in that case. Second, we have assumed that $C_{i}(x_{\neg i})$ contains some nonempty open set of feasible solutions. These two facts satisfy Prop. 4.2 of (Dutang, 2013), which guarantees lower and upper semicontinuity of $C_{i}$ in $x_{\neg i}$
@@ 2.3. As mixed complementarity problems
Since the optimization problems in (2) are all linear in the decision variable $x_{i}$, tensor games with tensor constraints can exactly be reformulated as mixed complementarity problems by way of the KKT conditions; we use this approach here to solve them. Note that in a bimatrix game, the corresponding complementarity problem is linear, but in this is not necessarily true in $n$-player tensor games, which have multilinear terms. The full mixed complementarity problem is as follows:
(5) $\nabla_{x_{i}}L_{i}(x,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty,$
@@
$Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}$ $\perp 0\leq\gamma_{i,j}\leq\infty,\ \forall j\in\mathcal{O}_{i}$
@@
(6) $L_{i}(x,\lambda,\gamma)=A_{i}\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{i,j}(Q_{j}\;[\cdot]\;x-\epsilon_{j})$
for all $i\in 1..n$ and $j\in\mathcal{O}_{i}$. The complementarity notation $F_{i}(Z)\perp l_{i}\leq Z_{i}\leq u_{i}$ means either $F_{i}(Z)\geq 0$ and $Z_{i}=l_{i}$, or $F_{i}(Z)=0$ and $l_{i}\leq Z_{i}\leq u_{i}$, or $F_{i}(Z)\leq 0$ and $Z_{i}=u_{i}$. The second complementarity condition in Eq. 5 is simply an equality ensuring that the mixing weights sum to 1; the corresponding dual $\lambda_{i}$ is free. Here, $\lambda_{i}$ are duals for the (implied) constraint that probabilities sum to 1, and $\gamma_{i,j}$ are duals for the tensor constraints.
3. Discrete Multiple-Player Constraints
With our necessary preliminaries complete, but before discussing pure strategy optimization, we pause to conceptualize tensor constraints as chance constraints to account for the probabilistic nature of mixing.
3.1. Chance Constrained Tensor Games
3.1.1. Definition

### v2

@@ Abstract.
When modeling robot interactions as Nash equilibrium problems, it is desirable to place coupled constraints which restrict these interactions to be safe and acceptable (for instance, to avoid collisions). Such games are continuous with potential mixed strategy equilibria, and this combination of characteristics means special care must be given to setting coupled constraints in a way that respects mixed strategies while remaining compatible with continuous game solution methods. Here, we investigate the problem of constraint-setting in this context, primarily focusing on a chance-based method. We first motivate these chance constraints in a discrete setting, placing them on $n$-player matrix games as a justifiable approach to handling the probabilistic nature of mixing. Then, we describe a numerical solution method for these chance constrained, continuous games with simultaneous pure strategy optimization. Finally, using a modified pursuit-evasion game as a motivating example, we demonstrate the actual behavior of this solution method in terms of its fidelity, parameter sensitivity, and efficiency.
Key words and phrases:
Game theory, mixed strategies, constrained optimization
@@ 1. Introduction
When designing models of multi-agent interactions in robotics and control, it is frequently necessary to place coupled constraints on agent behavior — for instance, preventing agents from colliding or jointly disrupting the environment. Competitive control scenarios are often described as continuous normal form games, but when coupled constraints are present, a choice is required about how to implement the constraints when many potential strategy profiles might be realized. In this work, we discuss proper engineering of constraints in a mixed-strategy, continuous framework, and put forth a chance-constraint-based approach as a compelling method.
We begin by discussing the behavior of coupled constraints in discrete games with mixed strategies. Since mixed strategies are probabilistic, a constraint may or may not be violated depending on the strategies actually realized by all players, and we describe the main intuitive methods used to account for this. Next, we consider continuous games. In general, continuous games require mixing over an infinite support of pure strategies to ensure the existence of a Nash equilibrium, but in certain classes of games such as separable games, finite support mixed strategy equilibria exist (Stein et al. (2008)), and constraints can be placed across these finite strategies. However, solver approaches for continuous games require the constraints to be differentiable, introducing another consideration Peters et al. (2022). To that end, we develop an iterative-tightening solver approach allowing coupled chance constraints to be used in continuous games.
To provide a concrete example, we consider a constrained, continuous, three-player pursuit-evasion game. We demonstrate the behavior of the chance-based constraint-setting approach (and an alternative), and analyze efficacy in maintaining the constraint, permissiveness of mixed strategy reasoning, and potential unintended consequences of the modeling choice. We conclude with promising areas of further investigation, and a discussion of the application of these constraints to model common problems in our spatial domain.
@@ 1.1. Our Contributions
  • A formulation of chance constraints applicable to mixed-strategy “tensor games with tensor constraints” (normal form games with coupled constraints);
  • A numerical solver which uses this formulation to find mixed-strategy Nash equilibria in continuous games with an iterative-tightening approach; and
  • Empirical behavior of this modeling approach (and an alternative) in a constrained pursuit-evasion game, including analysis of quality with respect to its parameters and the feasibility of the resulting solution.
The structure of this paper is as follows: In the remainder of this section we briefly discuss the related work on generalized Nash equilibrium problems and chance constraints to scope our work. In Section 2 we discuss “tensor games with tensor constraints” — a convenient set of normal form games with coupled constraints — and potential constraint-setting approaches. In Section 4, we move to the continuous setting and describe a solver with which we compute the empirical results discussed in Section 5. We close with a brief consideration of practical considerations and future work.
@@ 1.2.1. Coupled constraints
In computational game theory, coupled constraints are constraints which involve the decision variables of multiple players. These are particularly important in multi-agent robotics and control, where they prevent collisions or other undesirable interactions. Games which feature them are known as generalized Nash equilibrium problems, or GNEPs Facchinei and Kanzow (2010a). Older terminology refers to coupled constraints as feasible set interactions, found in social equilibrium problems and pseudo-Nash games Harker (1991).
Solutions to solving GNEPs vary, and do not typically use an explicit concept of mixed strategies. Krawczyk (2007) summarizes numerical solutions to solving these games, a project further advanced in, e.g., von Heusinger (2009). Past avenues include penalty methods Facchinei and Kanzow (2010b), augmented Lagrangian methods Kanzow and Steck (2016), and methods informed by error bound theory Ba and Pang (2022). This work focuses on the representation and modeling of constraints in mixed-strategy games, and less on computation. In this work we characterize GNEPs as a mixed complementarity problem via the KKT conditions, and solve them via a readily available solver Dirkse and Ferris (1995).
We diverge somewhat from this body of work in that we do explicitly consider mixed strategies, optimizing both pure strategies and mixing weights to solve continuous games. To our knowledge, constraint design in this form of generalized Nash equilibrium has not yet been fully explored, which prevents further development of mixed-strategy agents in robotics and control.
@@ 1.2.2. Chance constraints
To intuitively account for the stochasticity of mixed strategies, we use chance constraints, originally formulated in Charnes and Cooper (1959). Chance-constrained problems without feasible set interactions (that is, in ordinary Nash equilibrium problems) are well covered, and significant work has been done to characterize the circumstances in which such problems have equilibria (as in Henrion and Strugarek (2008), Geletu et al. (2013)). Chance constraints have been used in the presence of mixing for decades, see e.g. Mukherjee (1980).
Some work has been done describing the behavior of chance constraints in GNEPs. Zhang et al. (2022) manage mixing chance constraints in discrete mixed-strategy scenarios from the variational inequality point of view, while Singh and Lisser (2018) apply the same treatment to a continuous single strategy to handle stochastic payoffs, but the former assumes preselected pure strategies, while the latter does not mix. We assume pure strategy parameters are additional continuous decision variables rather than as given, and mix over them under the presence of chance constraints.
2. Formulation
In this section, we discuss discrete Nash equilibrium problems with coupled constraints — by definition, existence of equilibria, and potential constraint-setting approaches.
2.1. Tensor games with tensor constraints
We begin by describing a set of “generalized” normal form games (generalized in the sense of “generalized Nash equilibrium problems”; that is, normal form games with coupled constraints). This is not a novel formulation, but it is a convenient one for our purposes.
We work with normal form games as expressed with tensors, sometimes referred to as tensor games in computational game theory. Each player $i$ mixes over $m_{i}$ pure strategies with probabilities $x_{i}$. We denote player $i$’s $k$-th pure strategy $s_{i,k}$ and that player’s pure strategy space as $\mathcal{S}_{i}$, such that $s_{i}\in\mathcal{S}_{i}^{m_{i}}$ (and similarly denoting $s\in\mathcal{S}$).
Each player has a cost tensor, and each element of the cost tensor corresponds to a joint strategy. Cost tensors are denoted $A_{i}$, and all $A_{i}\in\mathbb{R}^{m_{1},..,m_{i}}$. The resulting cost, when each player has probability $x_{i,k}$ assigned to their $k$th strategy, is a homogeneous $N$-polynomial. Lacking more familiar notation, we denote the operation that forms this polynomial $A_{i}\;[\cdot]\;x$, under the following definition:
[…]
@@ 2.1. Tensor games with tensor constraints
$T\;[\cdot]\;x:=(((T\cdot x_{i})\cdot..)\cdot x_{N}$ (1)
We also constrain these games with a total of $C$ constraints (indexed by $j)$. Each player may or may not respect a particular constraint; we denote the index set of constraints respected by player $i$ as $\mathcal{O}_{i}$. Each constraint tensor is denoted as $Q_{j}$ and has the same shape as the cost tensor. The elements of the constraint tensor represent some value regarding the feasibility of the joint strategy with the corresponding index, and we constrain the expected value: $Q_{j}\;[\cdot]\;x\geq\epsilon_{j}$ for some arbitrary $\epsilon_{j}$.
With this notation in place, we define “tensor games with tensor constraints”:
[…]
@@
$\forall i\in\{1..N\}:\begin{bmatrix}\;\;\min\limits_{x_{i}\in\Delta^{m_{i-1}}}\;\;A_{i}\;[\cdot]\;x\;\;\;\;\textrm{s.t.}\;\;\;\;\\[8.0pt] \;\;\;\;\;\;Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}\geq 0\;\;\forall j\in\mathcal{O}_{i}\end{bmatrix}$ (2)
Let $x_{\neg i}$ denote the concatenation of strategy vectors
$x_{1},...,x_{i-1},x_{i+1},...,x_{N}$
and let $(x_{i},x_{\neg i})$ be shorthand for $x$. To be entirely explicit, we define Nash equilibrium on tensor games with tensor constraints:
[…]
A Nash equilibrium for a tensor game with tensor constraints $G(A,Q,\epsilon)$ is a strategy vector $x^{*}$ such that for all $i\in\{1,...,N\}$,
$\begin{split}A_{i}\;[\cdot]\;(x_{i}^{*},x_{\neg i}^{*})\leq A_{i}\;[\cdot]\;(x_{i},x_{\neg i}^{*})\\ \forall x_{i}\in\Delta^{m_{i}-1}:Q_{i,j}\;[\cdot]\;(x_{i},x_{\neg i}^{*})-\epsilon_{i,j}\geq 0\;\;\forall j\in\mathcal{O}_{i}\end{split}$ (3)
In discrete games, the way in which the contents of the cost and constraint tensors are derived is external to the game. However, to model continuous games, we assume that the elements of the tensors are related to a set pure strategies $s$ by way of real-valued cost and constraint functions $f_{i}$ and $g_{i,j}:\mathcal{S}\rightarrow\mathbb{R}$. (We implicitly assume players also choose a finite number of pure strategies from $\mathcal{S}$ which may or may not support a Nash equilibrium. We discuss this approach in greater detail Section 3.)
Let $K$ be an index into a cost or constraint tensor (and therefore $K_{i}$ is the index of the strategy realized by player $i$ in a particular joint strategy). Then, the cost and constraint tensor elements are defined accordingly:
$A_{i,K}=f_{i}(s_{1,K_{1}}\;..\;s_{N,K_{N}})\;\;;\;\;Q_{i,j,K}=\rho(g_{i,j}(s_{1,K_{1}}\;..\;s_{N,K_{N}}))$ (4)
@@
When solving a continuous tensor game with tensor constraints, our goal (from the perspective of player $i$) is to both mix over pure strategies correctly, and to select $s_{i}$ that generate the most favorable $A$ and $Q$ under a correct mix. For this purpose we assume $f$ and $g$ are twice differentiable and defined over a compact $\mathcal{S}$.
@@ 2.2. Existence of equilibria
Tensor games with tensor constraints do not necessarily admit a Nash equilibrium. As such, we briefly pause to establish sufficient conditions to ensure the presence of an equilibrium in a tensor-constrained tensor game.
{theorem}
@@
We then follow the conditions put forth by Dutang (2013):
@@
    First, $Q_{j}\;[\cdot]\;x$ is continuous and concave in $x_{i}$ given all $x_{\neg i}$, since it is linear in that case. Second, we have assumed that $C_{i}(x_{\neg i})$ contains some nonempty open set of feasible solutions. These two facts satisfy Prop. 4.2 of Dutang (2013), which guarantees lower and upper semicontinuity of $C_{i}$ in $x_{\neg i}$
@@ 2.3. As mixed complementarity problems
Tensor games with tensor constraints can exactly be reformulated as mixed complementarity problems by way of the Karush-Kuhn-Tucker (KKT) conditions. We use that mainstream approach here to solve them. Note that in a bimatrix game, the corresponding complementarity problem is linear, but in this is not necessarily true in $n$-player tensor games, which have multilinear terms. The full mixed complementarity problem is as follows:
$\nabla_{x_{i}}L_{i}(x,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty,$ (5)
@@
$Q_{j}\;[\cdot]\;x-\epsilon_{j}$ $\perp 0\leq\gamma_{j}\leq\infty,\ \forall j\in\mathcal{O}_{i}$
@@
$L_{i}(x,\lambda,\gamma)=A_{i}\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{j}(Q_{j}\;[\cdot]\;x-\epsilon_{j})$ (6)
for all $i\in 1..n$ and $j\in\mathcal{O}_{i}$. The complementarity notation $F_{i}(Z)\perp l_{i}\leq Z_{i}\leq u_{i}$ means either $F_{i}(Z)\geq 0$ and $Z_{i}=l_{i}$, or $F_{i}(Z)=0$ and $l_{i}\leq Z_{i}\leq u_{i}$, or $F_{i}(Z)\leq 0$ and $Z_{i}=u_{i}$. The second complementarity condition in Eq. 5 is simply an equality ensuring that the mixing weights sum to 1; the corresponding dual $\lambda_{i}$ is free. $\gamma_{i,j}$ are duals for the tensor constraints.
2.4. Using chance constraints in tensor games
We have yet to discuss $\rho_{j}$, the “constraint indicating function,” which we used to formalize coupled constraints in Section 2.1. Depending on the choice of $\rho_{j}$, constraints take on a different interpretation and different behavior. We conceptualize two possible choices in this and the next section.